how many gallons of a 70% acid solution must be mixed with 30 gallons of a 24% solution to obtain a solution that is 60% acid?
Answer by checkley77(12569)

.70x+.24*30=.60(30+x)
.70x+7.2=18+.6x
.70x-.60x=18-7.2
.1x=10.8
x=10.8/.1
x=108 gallons of 70% acid is used.
Proof:
.70*108+.24*30=.60(30+108)
75.6+7.2=.6*138
82.8=82.8